• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 19:24:45 on July 23, 2006

    Entry number 175065

    keyword=Retiming check

    I checked the cables for the retiming more upstream, two lemo cables looked suspicious so I replaced them, I also plugged the retiming signal in channel 4 of the EDTM scaler to check it.

    I also replugged T9 and fixed the end of run script to take it into account.